Operations Report – January 30th

Operational Report

Crew 188: Zac Trolley



This morning I was on an EVA, and it was cloudy. Due to the concern from CapCom about the battery bank dropping below 80% I decided to keep the generator running for the morning. Upon my return the generator was no longer running, so I do not have proper metrics for the generator. Shannon suggested that the car battery used to start the generator had died. Upon closer inspection, it the gauge for the battery charger is not working, and It’s suspect that the charger is working at all. I’ve swapped to the backup battery to run the generator tonight.

There as a new battery charger delivered, and it required maintenance to become operational (See attached picture) It is now charging the battery that was hooked up to the generator that died.

· Generator (hours run): Unknown

· Generator turned off: Unknown

· Generator turned on: 17:06 @ SOC 57%

Fuel Levels:

· Diesel: 82%

· Propane: 64%

· Ethanol Free Gasoline (5 Gallon containers for ATV): 8 Gallons

Water Levels:

· Water (trailer): 0 Gallons

· Water (static): 350 Gallons

· Trailer to Static Pump used: No

· Water (loft) – Static to Loft Pump used: Yes

· Meter @ 19:24 -130550.9

· Water Used between 7:00-19:00 37 Gal

· Toilet tank emptied: Yes


ATVs were not used today.

· ATVs Used: None

· Oil Added: None

· ATV Fuel Used: None


Spirit and Curiosity were used for today’s EVA. It was noted by the Director that Curiosity was not in the “off” position and not properly charging. This was crew oversight and likely caused by lower visibility in the suit helmets and not being able to see the charging light.

The Hab Director mentioned that the Rovers required an extension cord that is shorter than 25 feet. We estimate the cords being used now are 50 foot cables. I was able to find a 25 foot extension cable, meaning we could charge one rover at time based on the above specifications. Can we have the appropriate extension cables sent to the Hab in order to charge the Rovers?

Deimos rover used: No

· Hours: 106.9h

Spirit rover used: Yes

· Hours: 17.5h

· Beginning charge: 63%

Opportunity rover used: No

· Hours: 13.5h

· Beginning charge: 100%

Curiosity rover used: Yes

· Hours: 9.3h

· Beginning charge: 56%

HabCar: Not used

Summary of internet: We will be purchasing 5 gigs of data per day in order to meet our Sim objectives and be able to increase the visibility of this mission.

Summary of suits and radios: All nominal

Summary of Hab operations: All nominal

Summary of EVA operations: The helmets fogging continue to be an issue. We are going to bring the helmets into the open area in an effort warm up the Lexan before we go on an EVA. Please advise if there are other tricks and tip we can use to lower the fogging issue. It’s becoming a safety issue as visibility is now an unknow variable.

Summary of GreenHab operations: The temperature gauge display has some non-functional LCD areas, a new one has been ordered.

Summary of Science Dome operations: The science dome is awaiting a portable heater as per the Hab Director.

Summary of RAM operations: Not Operational.

Summary of health and safety issues: There was no report of a propane smell today, still awaiting the propane detector to rule out off gassing in the Hab.

Questions, concerns and requests to Mission Support: A crew member had previous version of the Hab hand book (2003, 2016) and those manuals mentioned higher functionality in the Hab in some areas (HALpr, weather station, CapCom and local chat room, etc) I am interested in bringing similar capability and more to the Hab; who is a good contact Earth side to discuss these improvements while I’m at the Hab and can investigate?

Copyright © The Mars Society. All rights reserved. | Main Site